OpenStudy (anonymous):

A plumber and an assistant finished a job in 4 hours, The job would have taken the plumber 6 h working alone. How long would the job have taken the assistant working alone?

OpenStudy (eyust707):

Is this for math or econ?

OpenStudy (anonymous):

math

OpenStudy (eyust707):

how many more hours would it have taken just the plumber?

OpenStudy (anonymous):

plumbers rate is \[\frac{1}{6}\] per hour combined rate is \[\frac{1}{4}\] per hour so assistants rate is \[\frac{1}{4}-\frac{1}{6}\] per hour

OpenStudy (anonymous):

the work, please.